### v1

# Addressing Confounding and Continuous Exposure Measurement Error Using Corrected Score Functions

Abstract
Confounding and exposure measurement error can introduce bias when drawing inference about the marginal effect of an exposure on an outcome of interest. While there are broad methodologies for addressing each source of bias individually, confounding and exposure measurement error frequently co-occur and there is a need for methods that address them simultaneously. In this paper, corrected score methods are derived under classical additive measurement error to draw inference about marginal exposure effects using only measured variables. Three estimators are proposed based on g-formula, inverse probability weighting, and doubly-robust estimation techniques. The estimators are shown to be consistent and asymptotically normal, and the doubly-robust estimator is shown to exhibit its namesake property. The methods, which are implemented in the R package mismex, perform well in finite samples under both confounding and measurement error as demonstrated by simulation studies. The proposed doubly-robust estimator is applied to study the effects of two biomarkers on HIV-1 infection using data from the HVTN 505 preventative vaccine trial.
Keywords Causal inference; Confounding; Corrected score functions; HIV/AIDS; Measurement error.
1 Introduction
Consider estimating the effect of a continuous exposure on an outcome when (i) the exposure is measured with error and (ii) the exposure-outcome association is potentially confounded. A motivating example is provided by the HVTN 505 trial of a preventive HIV vaccine. This HIV vaccine efficacy trial stopped administering immunizations early after reaching predetermined cutoffs for efficacy futility (Hammer et al., 2013). However, subsequent analyses of trial data identified several immunologic biomarker correlates of HIV acquisition among HIV vaccine recipients (Janes et al., 2017; Fong et al., 2018; Neidich et al., 2019). Some of these biomarkers could be possible target immune responses for future vaccines, so it is of interest to assess the effect of these biomarkers on risk of HIV acquisition. These biomarker effects must be assessed carefully since measurement of the biomarkers is subject to error and the association between the biomarkers and HIV risk is likely confounded.
Methods for estimating continuous exposure effects in the presence of confounding and exposure measurement error are limited. Existing measurement error methods can be used to target conditional treatment effects (i.e., the effects of a treatment on individuals conditional on their covariate values) by adjusting for baseline covariates in outcome models, for example, when performing regression calibration (Carroll et al., 2006). However, marginal effects (i.e., the average effects of a treatment across all individuals in a population) are often of primary interest. Josey et al. (2023) introduced a multiple imputation method that combines regression calibration and Bayesian approaches to estimate marginal causal effects in the presence of continuous exposure measurement error. The Josey et al. (2023) method requires replicate error-prone measurements from a cluster to estimate potential outcomes at cluster-level exposure values; moreover, it relies on a correctly specified outcome model. This paper develops methods which can be used when a continuous exposure is subject to measurement error and neither replicate measurements nor validation data are available. The approach is based on the corrected score (CS) function method for estimation and inference in the presence of measurement error (Carroll et al., 2006). Three CS methods targeting marginal causal estimands are proposed, based on the g-formula, inverse probability weighting (IPW), and doubly-robust (DR) estimation.
This paper proceeds as follows. In Section 2 notation and the target estimand are defined, and assumptions are stated. In Section 3 the three proposed estimators are introduced. In Section 4 the proposed estimators are evaluated in a simulation study, and in Section 5 one of the estimators is applied to study two biomarkers from the HVTN 505 vaccine trial. Section 6 concludes with a discussion of the advantages and limitations of the proposed methods.
2 Notation and Estimand
Suppose there are $m$ exposures/treatments of interest which may or may not be measured with error. Let $\boldsymbol{A}=(A_{1},A_{2},...,A_{m})$ be a row vector denoting the true exposure values for an individual and $\boldsymbol{A}^{*}=(A^{*}_{1},A^{*}_{2},...,A^{*}_{m})$ be the corresponding measured exposure values. For example, in the HVTN 505 trial, one biomarker of interest $A$ is antibody-dependent cellular phagocytosis activity. This biomarker was not observed exactly, but an imperfect phagocytic score $A^{*}$ was measured. Assume a classical additive measurement error model, i.e., that $\boldsymbol{A}^{*}=\boldsymbol{A}+\boldsymbol{\epsilon}_{me}$, where the measurement error $\boldsymbol{\epsilon}_{me}=(\epsilon_{me_{1}},\dots,\epsilon_{me_{m}})$ follows a multivariate normal distribution $\boldsymbol{\mathcal{N}}(\boldsymbol{0},\boldsymbol{\Sigma}_{me})$ with known covariance matrix $\boldsymbol{\Sigma}_{me}$. Without loss of generality, assume that only the first $j\leq m$ exposures are measured with error, such that $A_{k}=A^{*}_{k}$ for $k>j$ and the measurement error covariance has the block diagonal structure
$\boldsymbol{\Sigma}_{me}=\begin{bmatrix}\boldsymbol{\Sigma}_{e}&\boldsymbol{0}_{j,m-j}\\ \boldsymbol{0}_{m-j,j}&\boldsymbol{0}_{m-j,m-j}\end{bmatrix},$
where $\boldsymbol{\Sigma}_{e}$ is the measurement error covariance matrix for exposures $A_{1},...,A_{j}$ and $\boldsymbol{0}_{a,b}$ denotes an $a\times b$ matrix of zeros. The exposures may be either discrete or continuous, but all mismeasured exposures are assumed to be continuous.
Let $Y$ be the outcome of interest. Define $Y(\boldsymbol{a})$ to be the potential outcome under exposure $\boldsymbol{A}=\boldsymbol{a}=(a_{1},a_{2},...,a_{m})$. Assuming $j\geq 1$, there is at least one continuous exposure and each individual has an infinite number of potential outcomes. Let $\boldsymbol{L}=(L_{1},L_{2},...,L_{p})$ represent a vector of baseline covariates measured prior to exposure. This covariate set is assumed to be sufficient to adjust for confounding in the sense that the conditional exchangeability assumption described below is satisfied. Assume that $n$ i.i.d. copies of the random variables $(\boldsymbol{L},\boldsymbol{A}^{*},Y)$ are observed.
The estimand of interest is the mean dose-response surface, namely $\eta(\boldsymbol{a})\equiv\text{E}\{Y(\boldsymbol{a})\}$ for $\boldsymbol{a}\in\boldsymbol{\mathcal{A}}$, where $\boldsymbol{\mathcal{A}}$ represents the $m$-dimensional space of exposure values of interest. For example, with one exposure, $\eta(a)$ may be a dose response curve across an interval of exposure values. Each of the proposed estimators described in this paper will make assumptions that explicitly or implicitly impose restrictions on this surface. The proposed IPW estimator will directly target the parameters of a marginal structural model (MSM) given by
$\eta(\boldsymbol{a})=\eta(\boldsymbol{a};\boldsymbol{\gamma})$ (1)
where $\boldsymbol{\gamma}$ is a row vector quantifying the causal effects of the exposures on the outcome. For example, the MSM could be assumed to have the form $\eta(\boldsymbol{a};\boldsymbol{\gamma})=g^{-1}(\gamma_{0}+\boldsymbol{\gamma}_{a}\boldsymbol{a}^{T})$ for $\boldsymbol{\gamma}=(\gamma_{0},\boldsymbol{\gamma}_{a})$ and a monotone link function $g$ (e.g., the logit function).
The g-formula and doubly robust estimators, on the other hand, assume a parametric model for the mean outcome $\mu(\boldsymbol{L},\boldsymbol{A})\equiv\text{E}(Y|\boldsymbol{L},\boldsymbol{A})$ given the exposure and covariates:
$\mu(\boldsymbol{L},\boldsymbol{A})=\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta}),$ (2)
where $\boldsymbol{\beta}$ is row vector of parameters quantifying the association between $(\boldsymbol{L},\boldsymbol{A})$ and $Y$. For example, this outcome model could be assumed to have the form $\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})=g^{-1}(\beta_{0}+\bm{L}\boldsymbol{\beta}_{l}^{T}+\boldsymbol{A}\boldsymbol{\beta}_{a}^{T}+\bm{A}\boldsymbol{\beta}_{al}\boldsymbol{L}^{T})$, where $g$ is a monotone link function, $\boldsymbol{\beta}_{al}$ is an $m\times p$ matrix of interaction parameters with appropriate elements constrained to zero to include only relevant interactions, and $\boldsymbol{\beta}=[\beta_{0},\boldsymbol{\beta}_{l},\boldsymbol{\beta}_{a},\text{vec}(\boldsymbol{\beta}_{al})]$. The assumed mean outcome model implies the marginal structural model $\eta(\boldsymbol{a})=\text{E}\{\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta})\}$. The choice of specifying an MSM either explicitly with $\eta(\boldsymbol{a};\boldsymbol{\gamma})$ or implicitly with $\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})$ is a key consideration when choosing between methods.
The proposed methods in Section 3 rely on a standard set of assumptions used in causal inference: (i) causal consistency, $Y=Y(\boldsymbol{a})$ when $\boldsymbol{A}=\boldsymbol{a}$; (ii) conditional exchangeability, $Y(\boldsymbol{a})\perp\!\!\!\perp\boldsymbol{A}|\boldsymbol{L}$ for all $\boldsymbol{a}\in\boldsymbol{\mathcal{A}}$; and (iii) positivity, ${\color[rgb]{0,0,0}f_{\boldsymbol{A}|\boldsymbol{L}}}(\boldsymbol{a}|\boldsymbol{l})>0$ for all $\boldsymbol{l}$ such that ${\color[rgb]{0,0,0}f_{\boldsymbol{L}}}(\boldsymbol{l})>0$ and for all $\boldsymbol{a}\in\boldsymbol{\mathcal{A}}$. Here and below, the notation $f_{X}(x)$ denotes the marginal probability density function (pdf) of a random variable $X$, and $f_{X|W}(x|w)$ denotes the conditional pdf of $X$ given the random variable $W$. In addition, assume that the outcome and covariates are not measured with error, and that there is no model mis-specification unless otherwise stated.
Figure 1: An example directed acyclic graph (DAG), with variables defined in Section 2 and $\epsilon_{me_{1}}$ and $\epsilon_{me_{2}}$ corresponding to measurement error. This DAG represents a scenario with $m=4$ exposures, one each of the following: mismeasured and unconfounded ($A_{1}$), mismeasured and confounded ($A_{2}$), correctly measured and unconfounded ($A_{3}$), and correctly measured and confounded ($A_{4}$). In this case, $A_{3}=A^{*}_{3}$ and $A_{4}=A^{*}_{4}$ since they are both measured without error.
The estimators proposed in this paper are applicable in settings such as the example directed acyclic graph (DAG) in Figure 1. The methods can accommodate all four types of exposures in the DAG: unconfounded and correctly measured, unconfounded and mismeasured, confounded and correctly measured, and confounded and mismeasured. Here and throughout, exposure measurement error is assumed to be independent of $Y,\boldsymbol{A},\boldsymbol{L}$. For example, in Figure 1, $\boldsymbol{\epsilon}_{me}\perp\!\!\!\perp Y,\boldsymbol{A},\boldsymbol{L}$ since $A^{*}_{1}$ and $A^{*}_{2}$ are colliders on all paths between $\epsilon_{me_{1}},{\color[rgb]{0,0,0}\epsilon_{me_{2}}}$ and $Y,\boldsymbol{A},\boldsymbol{L}$.
3 Methods
The proposed estimators combine existing methods to (i) adjust for confounding using g-formula, inverse probability weighting, and doubly-robust techniques and (ii) correct exposure measurement error using CS functions. To begin, the CS method is briefly reviewed.
3.1 Review of Corrected Score Functions
Consider an M-estimator $\widehat{\boldsymbol{\theta}}_{0}$ of a parameter $\boldsymbol{\theta}$ in the absence of measurement error. That is, $\widehat{\boldsymbol{\theta}}_{0}$ solves the estimating equation
$\sum_{i=1}^{n}\boldsymbol{\Psi}_{0}(Y_{i},\boldsymbol{L}_{i},\boldsymbol{A}_{i},\boldsymbol{\theta})=\boldsymbol{0},$ (3)
for some score (or estimating) function $\boldsymbol{\Psi}_{0}$ that is unbiased, i.e., $\text{E}\{\boldsymbol{\Psi}_{0}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\theta})\}=\boldsymbol{0}$. Note that the traditional parametric IPW, g-formula, and doubly-robust estimators (ignoring measurement error) can all be expressed as M-estimators. Given that the true exposure values $\boldsymbol{A}$ are not observed, solving (3) is generally not possible. However, the observed data can in some cases be used to construct a corrected score function $\boldsymbol{\Psi}_{CS}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta})$ with the property that
$\text{E}\left\{\boldsymbol{\Psi}_{CS}\left(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta}\right)|Y,\boldsymbol{L},\boldsymbol{A}\right\}=\boldsymbol{\Psi}_{0}\left(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\theta}\right).$ (4)
By the law of iterated expectations, $\boldsymbol{\Psi}_{CS}$ is then an unbiased estimating function, and therefore the root $\widehat{\boldsymbol{\theta}}$ of $\sum_{i=1}^{n}\boldsymbol{\Psi}_{CS}(Y_{i},\boldsymbol{L}_{i},\boldsymbol{A}^{*}_{i},\boldsymbol{\theta})$ is an M-estimator for $\boldsymbol{\theta}$.
Novick and Stefanski (2002) provide one strategy to construct corrected score functions using Monte-Carlo simulation of complex variables, which can be applied to any estimating function $\boldsymbol{\Psi}_{0}$ that is conditionally unbiased, i.e., $\text{E}\{\boldsymbol{\Psi}_{0}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\theta})|\boldsymbol{A}\}=\boldsymbol{0}$. The method entails adding imaginary (i.e., complex-valued) measurement error to the observed $\boldsymbol{A}^{*}$, then taking a conditional expectation, given the observed data, of the real component of the complex-valued estimating function. The result is a function $\boldsymbol{\Psi}_{CS}$ of the observed data that satisfies (4). More formally, let
$\boldsymbol{\Psi}_{CS}\left(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta}\right)=\text{E}\left[\text{Re}\left\{\boldsymbol{\Psi}_{0}(Y,\boldsymbol{L},\widetilde{\boldsymbol{A}},\boldsymbol{\theta})\right\}|Y,\boldsymbol{L},\boldsymbol{A}^{*}\right],$ (5)
where $\widetilde{\boldsymbol{A}}=\boldsymbol{A}^{*}+i\widetilde{\boldsymbol{\epsilon}}$, $i=\sqrt{-1}$, $\text{Re}(\cdot)$ denotes the real component of a complex number, and $\widetilde{\boldsymbol{\epsilon}}\sim\boldsymbol{\mathcal{N}}(\boldsymbol{0},\boldsymbol{\Sigma}_{me})$ has the same distribution as the real measurement error $\boldsymbol{\epsilon}_{me}$ and is independent of all other variables. Note $\boldsymbol{\Psi}_{CS}$ is implicitly also a function of $\boldsymbol{\Sigma}_{me}$, which is assumed known. In some cases, the expectation in (5) has a closed form expression (see Section 3.5). In all cases, it can be approximated stochastically with i.i.d. simulated copies $\widetilde{\boldsymbol{\epsilon}}_{1},\dots,\widetilde{\boldsymbol{\epsilon}}_{B}$ of $\widetilde{\boldsymbol{\epsilon}}$. The Monte-Carlo corrected score (MCCS) function using $B$ replicates is given by
$\boldsymbol{\Psi}_{MCCS}^{B}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta})=B^{-1}\sum_{b=1}^{B}\text{Re}\left\{\boldsymbol{\Psi}_{0}(Y,\boldsymbol{L},\widetilde{\boldsymbol{A}}_{b},\boldsymbol{\theta})\right\},$ (6)
where $\widetilde{\boldsymbol{A}}_{b}=\boldsymbol{A}^{*}+i\widetilde{\boldsymbol{\epsilon}}_{b}$. By the weak law of large numbers (conditional on $Y,\boldsymbol{L},\boldsymbol{A}^{*}$),
$\boldsymbol{\Psi}_{MCCS}^{B}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta})$ converges in probability to $\boldsymbol{\Psi}_{CS}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta})$ as $B\to\infty$, so $\boldsymbol{\Psi}_{MCCS}^{B}$ can approximate $\boldsymbol{\Psi}_{CS}$ for a large number of replicates $B$. Note however that for any number of replicates $B\geq 1$, $\boldsymbol{\Psi}_{MCCS}^{B}$ satisfies (4), and therefore the root of $\sum_{i=1}^{n}\boldsymbol{\Psi}_{MCCS}^{B}(Y_{i},\boldsymbol{L}_{i},\boldsymbol{A}^{*}_{i},\boldsymbol{\theta})$ is an M-estimator.
3.2 G-formula CS Estimator
The first proposed method applies the CS framework to a traditional g-formula estimator. When there is no measurement error, the g-formula estimator entails fitting the outcome model in (2), then using the resulting parameter estimator $\widehat{\boldsymbol{\beta}}$ to estimate the dose-response curve as $\widehat{\eta}(\boldsymbol{a})=n^{-1}\sum_{i=1}^{n}\mu(\boldsymbol{L}_{i},\boldsymbol{a};\widehat{\boldsymbol{\beta}})$. This can be expressed as an M-estimator with estimating function
$\boldsymbol{\Psi}_{GF-0}(Y,\bm{L},\boldsymbol{A},\boldsymbol{\theta}_{GF})=\begin{bmatrix}\left\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\right\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\\ \eta(\boldsymbol{a})-\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta})\end{bmatrix},$ (7)
where $\boldsymbol{\theta}_{GF}=[\boldsymbol{\beta},\eta(\boldsymbol{a})]$, and $\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})$ is the partial derivative of $\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})$ with respect to $\boldsymbol{\beta}$. To estimate the dose response surface in practice, one can compute $\widehat{\eta}(\boldsymbol{a})$ for a large grid of points $\boldsymbol{a}$ in the space of interest. To accommodate exposure measurement error, the proposed g-formula CS function is $\boldsymbol{\Psi}_{GF-CS}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta}_{GF})=$
$\begin{bmatrix}\text{E}\left(\text{Re}\left[\left\{Y-\mu(\boldsymbol{L},\widetilde{\boldsymbol{A}};\boldsymbol{\beta})\right\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\widetilde{\boldsymbol{A}};\boldsymbol{\beta})\right]|Y,\boldsymbol{L},\boldsymbol{A}^{*}\right)\\ \eta(\boldsymbol{a})-\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta})\end{bmatrix},$ (8)
where $\widetilde{\boldsymbol{A}}$ is defined as in Section 3.1. Note that no correction is needed for the second row of $\boldsymbol{\Psi}_{GF-CS}$ since it does not involve $\boldsymbol{A}^{*}$.
3.3 IPW CS Estimator
Another common causal inference technique is to weight an estimator by the inverse probability/density of exposure(s) conditional on a set of covariates $\boldsymbol{L}$ that satisfy the conditional exchangeability assumption. IPW estimators can also be expressed as M-estimators and, in the absence of measurement error, the IPW estimating function is
$\boldsymbol{\Psi}_{IPW-0}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\theta}_{IPW})=\begin{bmatrix}\boldsymbol{\Psi}_{PS}(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})\\ SW(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})\left\{Y-\eta(\boldsymbol{A};\boldsymbol{\gamma})\right\}\partial_{\boldsymbol{\gamma}}\eta(\boldsymbol{A};\boldsymbol{\gamma})\end{bmatrix},$ (9)
where
$SW(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})=\frac{f_{\boldsymbol{A}}(\boldsymbol{A};\boldsymbol{\xi})}{f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L};\boldsymbol{\zeta})}$ (10)
are stabilized weights, $\boldsymbol{\Psi}_{PS}$ is an estimating function for fitting propensity score models (i.e., assumed models for $\boldsymbol{A}$ and $\boldsymbol{A}|\boldsymbol{L}$) with parameters $\boldsymbol{\xi}$ and $\boldsymbol{\zeta}$, $\boldsymbol{\theta}_{IPW}=(\boldsymbol{\gamma},\boldsymbol{\xi},\boldsymbol{\zeta})$, and $\boldsymbol{\gamma}$ are the MSM parameters from (1).
Heuristically, weighting works by creating a pseudo-population where confounding is no longer present. For continuous exposures, this is accomplished by weighting by the inverse of the joint density of exposures conditional on confounders. The unconditional joint density in the numerator of $SW$ is used to stabilize the weights, and does not need to be correctly specified in order for the estimator to be consistent. The second equation in (9) corresponds to a weighted estimator of the generalized linear regression of the outcome on only the exposure and is indexed by the MSM parameters of interest $\boldsymbol{\gamma}$ rather than the $\boldsymbol{\beta}$ parameters from equations (7) and (8). This regression model does not include $\boldsymbol{L}$ since weighting eliminates confounding.
Methods for fitting MSMs with multiple treatments often use weights of the form $SW=\prod_{j=1}^{m}{\color[rgb]{0,0,0}f_{A_{j}}}(A_{j})/{\color[rgb]{0,0,0}f_{A_{j}|\boldsymbol{L}}}(A_{j}|\boldsymbol{L})$, e.g., as in Hernán et al. (2001); to factorize the denominator in this way, the $m$ exposures $A_{1},\ldots,A_{m}$ are assumed to be independent conditional on $\boldsymbol{L}$. This assumption, which is made in the simulation study, allows one to avoid the data-intensive task of multivariate conditional density estimation, but it may be dubious in various applications, such as when a treatment has a direct effect on another treatment or when treatments have an unmeasured common cause. This assumption does have testable implications; see Zhang et al. (2012) for a related testing procedure. Alternatively, the denominator of $SW$ could be estimated using a mixed effects model (Tchetgen Tchetgen and VanderWeele, 2012).
When the weights $SW$ are known, such as in a conditionally randomized experiment, one does not need to fit models for $\boldsymbol{A}$ and $\boldsymbol{A}|\boldsymbol{L}$, and the estimating equation vector (9) can instead be solved with the first equation $\boldsymbol{\Psi}_{PS}$ removed and the true $SW$ used in the second equation. In observational studies, the weights are usually not known and must be estimated. Models used to estimate weights will be referred to as propensity score models; these can be fit using various previously described approaches. For simplicity, in the remainder of this paper, weight components corresponding to continuous exposures will be constructed using a ratio of normal densities estimated from linear models (Hirano and Imbens, 2004). To illustrate this, first consider the setting where the true exposures are observed. For continuous exposure $A_{j}$, a model of the form $A_{j}=\boldsymbol{\zeta}(1,\boldsymbol{L})^{T}+\epsilon_{ps}$ might be used, where $\epsilon_{ps}\sim\mathcal{N}(0,\sigma^{2}_{ps})$. Then based on the fitted model, the estimated conditional density ${\color[rgb]{0,0,0}f_{\boldsymbol{A}_{j}|\boldsymbol{L}}}(A_{j}|\boldsymbol{L};\widehat{\boldsymbol{\zeta}})$ is used to estimate ${\color[rgb]{0,0,0}f_{\boldsymbol{A}_{j}|\boldsymbol{L}}}(A_{j}|\boldsymbol{L})$. An intercept-only model is used similarly to estimate the weight numerators. Other methods and more flexible choices for weight models (Naimi et al., 2014) can also be used for continuous exposures in practice.
When a parametric propensity model is used for the denominator of $SW$, integrating the assumed $\boldsymbol{A}|\boldsymbol{L}$ model over the covariate distribution implies certain restrictions on the marginal distribution of $\boldsymbol{A}$. However, the consistency of the IPW estimator does not rely on a correctly specified model for the numerator of $SW$, such that the numerator and denominator models do not necessarily have to be compatible. Finally, weight components corresponding to discrete exposures (which are assumed to be always correctly measured) can be estimated using common approaches, such as logistic and multinomial regression.
To account for mismeasured exposures, the IPW CS function is
$\boldsymbol{\Psi}_{IPW-CS}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta}_{IPW})=$
$\begin{bmatrix}\widetilde{\boldsymbol{\Psi}}_{PS}(\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\xi},\boldsymbol{\zeta})\\ \text{E}\left(\text{Re}\left[SW(\boldsymbol{L},\widetilde{\boldsymbol{A}},\boldsymbol{\xi},\boldsymbol{\zeta})\left\{Y-\eta(\widetilde{\boldsymbol{A}};\boldsymbol{\gamma})\right\}\partial_{\boldsymbol{\gamma}}\eta(\widetilde{\boldsymbol{A}};\boldsymbol{\gamma})\right]|Y,\boldsymbol{L},\boldsymbol{A}^{*}\right)\end{bmatrix},$ (11)
where $\widetilde{\boldsymbol{\Psi}}_{PS}$ is an estimating function to fit the marginal and conditional densities $f_{\boldsymbol{A}}(\boldsymbol{A};\boldsymbol{\xi})$ and $f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L};\boldsymbol{\zeta})$ for the true exposure $\boldsymbol{A}$ using the measured $\boldsymbol{A}^{*}$. Note these propensity models have a mismeasured outcome and, assuming the measurement error is additive and $\boldsymbol{\Sigma}_{me}$ is known, can be fit using existing methods, e.g., regression calibration or simulation-extrapolation (Carroll et al., 2006). In the case where $\boldsymbol{A}|\boldsymbol{L}$ follows a multivariate normal distribution with conditionally independent outcomes and a linear mean model, the estimating functions corresponding to ordinary least squares (OLS) with $\boldsymbol{A}^{*}$ in place of $\boldsymbol{A}$ will give consistent estimators of the mean model coefficients, and the usual OLS covariance estimator is consistent for $\text{Cov}(\boldsymbol{A}^{*}|\boldsymbol{L})=\text{Cov}(\boldsymbol{A}|\boldsymbol{L})+\boldsymbol{\Sigma}_{me}$, from which $\text{Cov}(\boldsymbol{A}|\boldsymbol{L})$ can be recovered since $\boldsymbol{\Sigma}_{me}$ is assumed known.
3.4 Doubly-Robust CS Estimator
Both g-formula and IPW methods rely on model specifications that may not be correct in practice. The g-formula provides consistent estimation of potential outcome means only when the outcome model (2) conditional on exposures and confounders is correctly specified. Likewise, IPW estimators are consistent only when the propensity score models (and the MSM in (1)) are correctly specified. In contrast, doubly-robust (DR) estimators entail specifying both propensity and outcome models, but remain consistent if one model is mis-specified and the other is not (Robins, Rotnitzky, and Zhao, 1994; Lunceford and Davidian, 2004; Bang and Robins, 2005).
One such DR method is a weighted regression estimator adapted from Zhang et al. (2016) for the additive measurement error setting. In general, this method only applies when the outcome model and the implied MSM are both linear on the scale of the same link function. That is, when $\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})=g^{-1}(\beta_{0}+\bm{L}\boldsymbol{\beta}_{l}^{T}+\boldsymbol{A}\boldsymbol{\beta}_{a}^{T}+\bm{A}\boldsymbol{\beta}_{al}\boldsymbol{L}^{T})$ and $\eta(\boldsymbol{a};\boldsymbol{\gamma})=g^{-1}(\gamma_{0}+\boldsymbol{\gamma}_{a}\boldsymbol{a}^{T})$ for the same $g$. The following two special cases fit into this class of models.
  • Case A: a linear outcome model (where $g$ is the identity link), which corresponds to a linear MSM with $\gamma_{0}=\beta_{0}+\text{E}(\boldsymbol{L})\boldsymbol{\beta}_{l}^{T}$ and $\boldsymbol{\gamma}_{a}=\boldsymbol{\beta}_{a}+\text{E}(\boldsymbol{L})\boldsymbol{\beta}_{al}^{T}$.
  • Case B: a log-linear outcome model (where $g$ is the $\log$ link) without $\boldsymbol{A}\times\boldsymbol{L}$ interactions (i.e., $\boldsymbol{\beta}_{al}=\boldsymbol{0}$), which corresponds to a log-linear MSM with $\gamma_{0}=\beta_{0}+\log[\text{E}\{\exp(\boldsymbol{L}\boldsymbol{\beta}_{l}^{T})\}]$ and $\boldsymbol{\gamma}_{a}=\boldsymbol{\beta}_{a}$.
The DR estimator is similar to the standard g-formula estimator, but with a weighted outcome regression where the weights are the inverse probability weights given in equation (10). With perfectly measured exposures, the DR estimator has corresponding estimating function $\psi_{DR-0}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\theta}_{DR})=$
$\begin{bmatrix}\boldsymbol{\Psi}_{PS}(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})\\ SW(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\\ \eta(\boldsymbol{a})-\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta})\end{bmatrix},$ (12)
where $\boldsymbol{\theta}_{DR}=(\boldsymbol{\theta}_{GF},\boldsymbol{\xi},\boldsymbol{\zeta})$. The corresponding DR CS estimating function is then
$\psi_{DR-CS}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\theta}_{DR})=$
$\begin{bmatrix}\widetilde{\boldsymbol{\Psi}}_{PS}(\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\xi},\boldsymbol{\zeta})\\ \text{E}\left(\text{Re}\left[SW(\boldsymbol{L},\widetilde{\boldsymbol{A}},\boldsymbol{\xi},\boldsymbol{\zeta})\{Y-\mu(\boldsymbol{L},\widetilde{\boldsymbol{A}};\boldsymbol{\beta})\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\widetilde{\boldsymbol{A}};\boldsymbol{\beta})\right]|Y,\boldsymbol{L},\boldsymbol{A}^{*}\right)\\ \eta(\boldsymbol{a})-\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta})\end{bmatrix}.$ (13)
If the stabilized weights $SW$ are unknown, they are estimated with the root of $\widetilde{\boldsymbol{\Psi}}_{PS}$ as described for the IPW CS estimator; if the weights are known, the estimating equation vector (13) can instead be solved with the first element $\widetilde{\boldsymbol{\Psi}}_{PS}$ removed. Like the g-formula CS estimator, the DR CS estimator can be evaluated over a grid of values $\boldsymbol{a}$ of interest to estimate a dose-response surface. Likewise, the doubly robust property relies on the MSM implied by the outcome model being correct. This condition is always met when the outcome model is correctly specified, but is not always met for an incorrect outcome model. For this reason, the doubly robust estimator is only recommended in scenarios where the specified outcome model is believed to be compatible with the true MSM.
3.5 Evaluating Corrected Score Functions
The three proposed CS methods involve expectations of the form $\text{E}[\text{Re}\{\boldsymbol{\Psi}_{0}(Y,\boldsymbol{L},\widetilde{\boldsymbol{A}},\boldsymbol{\theta})\}|Y,\boldsymbol{L},\boldsymbol{A}^{*}]$. The MCCS method described in Section 3.1 for approximating this expectation is convenient since it can be applied in a variety of modeling scenarios without dealing with complex algebra. This simulation-based approach can however be computationally intensive, and if this is a concern, there are some scenarios where a closed form expression for the CS function exists. Section 7.4.3 of Carroll et al. (2006) lists several such scenarios, including when $\boldsymbol{\Psi}_{0}$ is a score function for (i) normal linear regression with an identity link, (ii) Poisson regression with a log link, or (iii) gamma regression with a log link. If the outcome model (2) falls into any of these three cases, then the closed form expressions in Carroll et al. (2006) can be used for the g-formula estimator, bypassing any Monte-Carlo approximation.
If the weights $SW$ are known, then the IPW (or DR) CS function can be found in the same way for appropriate MSMs (or outcome models). However, when $SW$ are estimated, this strategy cannot be employed directly. In Web Appendix B, it is shown that in the special case where the models for $\boldsymbol{A}$ and $\boldsymbol{A}|\boldsymbol{L}$ are normal with identity links and where the MSM has a linear link, the IPW CS function has a closed form. Similarly, if the same holds for the $\boldsymbol{A}$ and $\boldsymbol{A}|\boldsymbol{L}$ models and if the outcome model has a linear link, then the DR CS function has a closed form expression.
3.6 Large-Sample Properties and Variance Estimation
In Web Appendix A, each of the three proposed estimators is shown to be consistent and asymptotically normal by showing that their corresponding estimating functions have expectation $\boldsymbol{0}$ (Stefanski and Boos, 2002). In addition, the DR CS estimator is shown to be consistent when only one of the propensity and outcome models is correctly specified.
Since each proposed method is an M-estimator, consistent estimators of their asymptotic variances are given by the empirical sandwich variance technique. Estimating equations corresponding to the estimation of weights should be included in the estimating equation vector for each method when computing the sandwich variance estimator. Wald $100(1-\alpha)\%$ confidence intervals (CIs) for the parameters of interest can then be constructed in the usual fashion. The sandwich variance estimator is consistent for the asymptotic variance but tends to underestimate the variance in finite samples, leading to undercoverage of the CI. To alleviate this, a bias-corrected (BC) sandwich variance estimator from Fay and Graubard (2001) can be used.
3.7 Handling Unknown Measurement Error Covariance
Although the proposed methods require no individual-level supplemental data, a priori knowledge of the measurement error covariance matrix is required. Sometimes this matrix will be known from properties of the measurement devices (e.g., some bioassays, certain widely studied survey instruments). Other times, however, $\boldsymbol{\Sigma}_{me}$ may need to be estimated or may be completely unknown. Here some guidelines are provided for such analyses.
In some cases, an estimate $\widehat{\boldsymbol{\Sigma}}_{me}$ of the measurement error covariance is available from a previous study; this can be viewed as summary-level supplemental data. In other cases, supplemental data from the current study may be available in the form of replicates of the potentially mismeasured variables. These replicates can be used to estimate $\boldsymbol{\Sigma}_{me}$ as described in Carroll et al. (2006). In particular, suppose for individual $i$ there are $k_{i}$ replicates of the mismeasured exposures, $\boldsymbol{A}^{*}_{i1},...,\boldsymbol{A}^{*}_{ik_{i}}$ with mean $\boldsymbol{A}^{*}_{i.}$. Then an estimator for the measurement error covariance is given by:
$\widehat{\boldsymbol{\Sigma}}_{me}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{k_{i}}(\boldsymbol{A}^{*}_{ij}-\boldsymbol{A}^{*}_{i.})^{T}(\boldsymbol{A}^{*}_{ij}-\boldsymbol{A}^{*}_{i.})}{\sum_{i=1}^{n}(k_{i}-1)}$
Likewise if validation data are available, the measurement error covariance can be estimated e.g., using maximum likelihood. There are many types of studies for which the covariance matrix may be assumed to follow a certain structure (e.g., diagonal, meaning the measurement errors are uncorrelated). For example, biological assays run on different samples and analyzed by separate machines/researchers may have uncorrelated or only weakly correlated measurement errors (Farrance and Frenkel, 2012). In these cases $\widehat{\boldsymbol{\Sigma}}_{me}$ can be modified to fit the assumed structure. For other types of data such as survey responses, analysts should be more cautious, noting that response bias, recall bias, and other forms of measurement error in survey instruments may be correlated within individuals (Biemer et al., 2013). Regardless of how the measurement error covariance is estimated, the methods described in this paper can be used with $\boldsymbol{\Sigma}_{me}$ replaced with $\widehat{\boldsymbol{\Sigma}}_{me}$. Section 7.5.2 of Carroll et al. (2006) provides one way to estimate the covariance of the the estimator $\widehat{\boldsymbol{\theta}}$, which accounts for uncertainty in $\widehat{\boldsymbol{\Sigma}}_{me}$. Alternatively, provided $\widehat{\boldsymbol{\Sigma}}_{me}$ is an M-estimator, the estimating equations for $\widehat{\boldsymbol{\Sigma}}_{me}$ and $\widehat{\boldsymbol{\theta}}$ can be stacked and solved simultaneously, and the corresponding sandwich variance estimator will reflect uncertainty in $\widehat{\boldsymbol{\Sigma}}_{me}$ (Cole et al., 2023; Shook‐Sa et al., 2024).
When no replicates are available and there is no prior knowledge of $\boldsymbol{\Sigma}_{me}$, the proposed methods can still be used in conjunction with a sensitivity analysis. In some settings, an upper bound on the exposure measurement error covariance may be assumed; for example, the covariance of a correctly measured exposure (if estimated in a prior study) may be a reasonable upper bound on measurement error covariance for the corresponding mismeasured exposure. Once upper bounds are determined, inference may be repeated using the proposed methods for a range of $\boldsymbol{\Sigma}_{me}$ specifications to assess robustness of point estimates and confidence intervals to the degree of assumed measurement error; this procedure is more straightforward when the matrix $\boldsymbol{\Sigma}_{me}$ is small and diagonal, and becomes difficult to interpret as the number of non-zero parameters grows.
4 Simulation Study
The performance of the proposed methods was evaluated in three simulation studies. The first simulation examined the proposed g-formula CS approach in a scenario where confounding and additive exposure measurement error were present. 1000 data sets of $n\in\{800,8000\}$ individuals were simulated, each with the following variables: a confounder $L_{1}$ simulated as a Bernoulli random variable with expectation $0.5$, i.e., $L_{1}\sim\text{Bern}(0.5)$, a second confounder $L_{2}\sim\text{Bern}(0.2)$, an exposure $A\sim\mathcal{N}(2+0.3L_{1}-0.5L_{2},0.6)$, and an outcome $Y\sim\text{Bern}(\text{logit}^{-1}(-2+0.7A-0.6L_{1}+0.4L_{2}-0.4AL_{1}-0.2AL_{2}))$. The exposure was subject to additive measurement error, simulated as $A^{*}\sim\mathcal{N}(A,0.25)$.
Dose-response curves were estimated using three methods: (i) oracle g-formula using the true values $A$ (which are unknown in practice), (ii) naive g-formula, which assumes $A^{*}$ are perfectly measured, and (iii) g-formula CS, the proposed estimator. Table 1 shows the empirical bias, empirical standard error (ESE), average estimated standard error (ASE), and percentage of confidence intervals (CIs) including the true value for the point on the dose response curve $\eta(3)$. ASE and CI coverage are shown for both the uncorrected (UC) and BC variance estimators.
[…]
UC BC
n Method Bias ESE ASE Cov ASE Cov
800 Oracle G-Formula 0.1 2.7 2.8 94.5 2.8 94.9
Naive G-Formula -2.8 2.4 2.4 76.4 2.5 76.9
G-Formula CS 0.6 4.9 3.2 92.0 5.2 94.4
8000 Oracle G-Formula 0.0 0.9 0.9 93.7 0.9 93.8
Naive G-Formula -2.9 0.8 0.8 3.2 0.8 3.2
G-Formula CS 0.0 1.0 1.0 94.8 1.0 94.8
The oracle g-formula estimator, which serves as a benchmark since it uses the unobserved $A$, was essentially unbiased, and its CIs achieved nominal coverage. The naive g-formula estimator had considerable bias, and its corresponding CIs had low coverage; this highlights the consequences of ignoring measurement error. The g-formula CS estimator on the other hand had small bias which decreased as $n$ increased, and its corresponding CIs had close to nominal coverage. The UC CI for the g-formula CS estimator with $n=800$ had some undercoverage (92%), which improved with either the BC CI or $n=8000$.
To demonstrate performance of the proposed estimator for the entire dose-response curve, empirical biases of the oracle, naive, and CS g-formula methods over a grid of points $a\in[1,4]$ and for $n=8000$ are displayed in Figure 2. The proposed g-formula CS estimator was approximately unbiased, while the naive g-formula method exhibited large bias across nearly the entire range of exposure values considered.
Figure 2: Estimated dose-response curve bias for the oracle, naive, and CS g-formula estimators in the first simulation study. Bias refers to the average bias across 1000 simulated data sets for each method evaluated at each point on the horizontal axis, corresponding to setting the true exposure to $a\in[1,4]$.
The second simulation study examined the proposed IPW CS approach. A total of 1000 data sets of $n\in\{800,8000\}$ individuals were simulated, each with the following variables: confounder $L\sim\text{Uniform}(0,1)$, exposure $\boldsymbol{A}=(A_{1},A_{2},A_{3})$ with $\boldsymbol{A}|L$ having multivariate normal distribution with mean $(0.4L,-0.4L,0.2-0.1L)$ and diagonal covariance matrix with diagonal elements $(0.09,0.09,0.09)$, and binary outcome $Y$ with $\text{E}(Y|L,\boldsymbol{A})=0.4+0.15A_{1}+0.15A_{2}+0.2A_{3}+0.1L+0.1LA_{1}-0.1LA_{3}$. This implied an MSM of $\eta(\boldsymbol{a};\boldsymbol{\gamma})=\gamma_{0}+\gamma_{1}a_{1}+\gamma_{2}a_{2}+\gamma_{3}a_{3}=0.45+0.2a_{1}+0.15a_{2}+0.15a_{3}$. The third exposure $A_{3}$ was correctly measured, while $A_{1}$ and $A_{2}$ were subject to additive measurement error simulated as $A_{1}^{*}\sim\mathcal{N}(A_{1},0.05)$ and $A_{2}^{*}\sim\mathcal{N}(A_{2},0.05)$ (i.e., the measurement error covariance matrix was diagonal).
The MSM parameter of interest $\boldsymbol{\gamma}=(\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3})$ was estimated for each simulated data set using three methods: (i) oracle IPW using the true values $A$, (ii) naive IPW treating the $A^{*}$ as perfectly measured, and (iii) the proposed IPW CS estimator. The empirical bias, ESE, ASE, and CI coverage are presented in Table 2.
Table 2: Results from the second simulation study. UC, BC, n, Bias, ASE, ESE, and Cov defined as in Table 1.
UC BC
n Method Parameter Bias ESE ASE Cov ASE Cov
800 Oracle IPW $\gamma_{0}$ 0.0 3.6 3.2 90.4 3.2 90.7
$\gamma_{1}$ 0.0 8.1 7.3 91.7 7.4 92.0
$\gamma_{2}$0.1 8.1 7.3 91.4 7.4 92.0
$\gamma_{3}$0.3 7.6 6.9 92.6 7.0 92.9
Naive IPW $\gamma_{0}$ 0.5 2.9 2.7 93.1 2.7 93.3
$\gamma_{1}$7.2 5.8 5.4 72.3 5.5 73.0
$\gamma_{2}$5.4 5.8 5.4 82.1 5.5 82.6
$\gamma_{3}$0.3 6.9 6.5 92.2 6.5 92.3
IPW CS $\gamma_{0}$ 0.0 4.8 4.0 91.0 4.1 91.7
$\gamma_{1}$ 0.1 12.1 10.4 89.3 10.7 90.5
$\gamma_{2}$0.2 12.2 10.4 89.9 10.7 91.3
$\gamma_{3}$0.5 8.7 7.7 91.1 7.8 91.2
8000 Oracle IPW $\gamma_{0}$ 0.0 1.2 1.1 95.1 1.1 95.1
$\gamma_{1}$ 0.1 2.8 2.7 94.1 2.7 94.3
$\gamma_{2}$ 0.0 2.8 2.6 93.6 2.6 93.8
$\gamma_{3}$ 0.1 2.4 2.4 94.5 2.4 94.5
Naive IPW $\gamma_{0}$ 0.5 0.9 0.9 91.6 0.9 91.6
$\gamma_{1}$7.1 1.9 1.8 2.5 1.8 2.6
$\gamma_{2}$5.4 1.9 1.8 16.1 1.8 16.1
$\gamma_{3}$ 0.1 2.2 2.1 94.7 2.1 94.7
IPW CS $\gamma_{0}$0.1 1.8 1.6 93.0 1.6 93.3
$\gamma_{1}$ 0.4 4.6 4.1 92.2 4.2 92.6
$\gamma_{2}$0.2 4.6 4.1 92.3 4.2 92.4
$\gamma_{3}$ 0.1 3.1 2.9 93.5 2.9 93.5
The oracle IPW estimator, which again serves as a benchmark, was unbiased and its CIs achieved nominal coverage. The naive IPW estimators of the intercept $\gamma_{0}$ and the coefficient $\gamma_{3}$ corresponding to the perfectly measured $A_{3}$ had low bias, but the naive estimators of $\gamma_{1}$ and $\gamma_{2}$ were biased and their CIs had low coverage. The IPW CS estimator on the other hand was essentially unbiased. The UC CI for the IPW CS estimator had some undercoverage, and the BC CI had closer to nominal coverage; this improved with increasing $n$.
The third simulation study compared the g-formula and IPW CS approaches to the DR CS estimator under various model specifications. In particular, 1000 datasets of $n=2000$ individuals were simulated with the following variables: confounders $L_{1}\sim\text{Binom}(0.5)$ and $L_{2}\sim\mathcal{N}(1,0.5)$, exposure $A\sim\mathcal{N}(2+0.9L_{1}-0.6L_{2},1.1)$, and continuous outcome $Y\sim\mathcal{N}(1.5+0.7A+0.9L_{1}-0.7AL_{1}-0.6L_{2}+0.4AL_{2},0.16)$ such that the assumptions of all three methods hold given correct model specifications. For this data generating process, the corresponding MSM was $\eta(a;\boldsymbol{\gamma})=\gamma_{0}+\gamma_{1}a$, where $\gamma_{0}=1.35$ and $\gamma_{1}=0.75$. The methods were compared with respect to their performance estimating $\gamma_{1}$. For the IPW estimator, this parameter was estimated directly. For the g-formula and DR estimators, this parameter was indirectly estimated as $\widehat{\eta}(1)-\widehat{\eta}(0)$; a corresponding variance estimator was constructed using the delta method. The exposure $A$ was subject to additive measurement error simulated as $A^{*}\sim\mathcal{N}(A,0.16)$. The three approaches were compared under scenarios where only the propensity model was correctly specified, only the outcome regression was correctly specified, both were correctly specified, or neither was correctly specified. The propensity model was mis-specified by not including the confounder $L_{1}$ and the outcome regression was mis-specified by leaving out $L_{1}$ and the interaction between $A$ and $L_{1}$.
Table 3: Results from the third simulation study. UC, BC, Bias, ASE, ESE, and Cov defined as in Table 1. PS indicates the propensity score model is correctly specified; OR indicates the outcome regression is correctly specified.
UC BC
Correct Specifications Method Bias ESE ASE Cov ASE Cov
PS and OR DR CS 0.5 2.3 2.1 91.8 2.1 92.2
G-Formula CS 0.1 1.6 1.6 94.8 1.6 94.8
IPW CS 0.2 7.3 4.7 91.9 5.2 93.2
PS DR CS 0.3 4.0 3.3 91.7 3.3 92.1
G-Formula CS -6.7 1.8 1.8 5.3 1.8 5.3
IPW CS 0.2 7.3 4.7 91.9 5.2 93.2
OR DR CS 0.2 1.8 1.7 94.4 1.7 94.3
G-Formula CS 0.1 1.6 1.6 94.8 1.6 94.8
IPW CS -6.5 2.4 2.2 17.4 2.2 18.0
DR CS -6.4 2.2 2.0 14.5 2.0 14.3
G-Formula CS -6.7 1.8 1.8 5.3 1.8 5.3
IPW CS -6.5 2.4 2.2 17.4 2.2 18.0
The simulation results are presented in Table 3 and support the theoretical results described in Section 3.4. Namely, when only the propensity score model was specified correctly, the IPW and DR CS estimators performed well, but the g-formula CS estimator was subject to substantial bias, and the corresponding CIs had lower than nominal coverage. Likewise when only the outcome model was specified correctly, the g-formula and DR CS estimators performed well, but the IPW CS estimator was biased, and the corresponding CIs had low coverage. That is, the DR estimator performed well when either one of the two models was mis-specified, exhibiting its namesake double-robustness property. When both models were mis-specified, the DR CS estimator performed similarly to the singly-robust approaches. In general, the BC CIs tended to have slightly better coverage than the UC CIs in settings where the corresponding estimator was consistent.
While all simulations in this section generated data under positivity and additive measurement error, simulations of the proposed methods under violations of the positivity and additive measurement error assumptions are presented in Web Appendix D. Finally, each of the three simulation studies were repeated under a sample size of $n=400$ to evaluate performance with smaller sample size. These results are also presented in Web Appendix D and yielded similar conclusions to the original simulations with larger sample size, except with larger finite-sample bias and undercoverage of the CIs.
5 Application
[…]
In this application, data are analyzed under a log link, i.e., a log-linear risk model, such that the proposed doubly-robust CS estimator can be used. Notably, the immunologic markers were not measured in all participants, but rather were measured in all vaccine recipients with HIV acquisition and in a stratified random sample of vaccine recipients who completed 24 months of follow-up without HIV acquisition. To account for this two-phase sampling design, the weights in the doubly-robust estimator are multiplied by inverse probability of sampling weights, following the procedure described in Wang et al. (2009). This version of the proposed estimator is described in more detail and evaluated in a simulation study in Web Appendix C. ADCP and RII were modeled separately to match the univariate-style analysis performed in Neidich et al. (2019); accordingly, separate propensity models were fit for each exposure. Thus, the results cannot be interpreted as the joint effect of the two exposures.
For each propensity model specification, main effects for the covariates age, race, BMI, behavior risk, CD4-P, and CD8-P were included. For the outcome model specification, main effects for the exposure of interest, age, race, BMI, behavior risk, CD4-P, and CD8-P were included. Inverse probability of sampling weights were computed based on the case-cohort sampling design, with different weights estimated for cases and non-cases. The weight denominators were estimated with sample proportions (i.e., for non-cases, using the proportion of non-cases who had immunologic markers measured at Month 7). Based on the theoretical and empirical results in Sections 3 and 4, the doubly-robust CS estimator should be consistent if either specification is correct. Model diagnostics for the propensity and outcome models are presented in Web Appendix E. Finally, each exposure was assumed to follow a classical additive measurement error model where a sensitivity analysis was performed by varying the measurement error variances within 0, 1/8, 3/16, and 1/4 times the variance for each exposure variable when restricted to vaccine recipients with an immunologic marker. Since ADCP and RII are log-transformations of strictly positive random variables, this setup is equivalent to assuming that their corresponding non-log transformed variables follow multiplicative measurement error models. All covariates were assumed to be measured without error. Dose response curves were estimated for each exposure and each measurement error level across a range of 0.5 to 3 for ADCP and 7 to 10 for RII; exposure levels above 3 and 10 respectively were associated with no or almost no risk.
Figure 3: HVTN 505 results. Each panel shows the dose-response curve for ADCP or RII estimated by the DR CS estimator, as well as their respective shaded 95% pointwise confidence regions. From top to bottom, each panel reflects increasing user-specified variances of measurement error $\sigma^{2}_{me}$ corresponding to proportions of 0, 1/8, 3/16, and 1/4 of $\sigma^{2}$, the estimated total exposure variances among vaccinees with immune responses.
The analysis results are plotted in Figure 3. For each exposure, lower values corresponded to higher HIV risk among the vaccine recipients, in line with prior results and biological theory. Moving across panels from top to bottom in Figure 3, the assumed measurement error variance increases and the confidence regions become wider, as expected. While higher levels of each exposure appear to cause lower risk, there is substantial uncertainty when allowing for higher magnitudes of measurement error and because of the low sample size (only 25 acquisition cases with measured Month 7 exposures). Thus, studies measuring these biomarkers in more participants are needed to draw stronger conclusions.
6 Discussion
In this paper, methods are proposed which adjust for both confounding and continuous exposure measurement error. The corrected score method, under an additive measurement error assumption, is applied to g-formula, IPW, and doubly-robust estimators. Empirical results confirm the results expected based on the large sample properties of the estimators. Accompanying this paper, an R package mismex has been developed.
While the proposed methods were shown to have favorable theoretical and empirical properties, they are not without limitation. In particular, the methods require that the measurement error covariance is known or can be estimated using data from previous studies or the current study. As demonstrated in Section 5, if the covariance is unknown then sensitivity analysis can be straightforward and informative if the covariance matrix is small or restricted such that it has few parameters. The proposed methods also rely on the additive measurement error assumption and correct specification of certain parametric models. The DR CS estimator provides some protection from mis-specification of outcome and propensity score models, requiring only one but not necessarily both to be correctly specified. When both models are mis-specified, DR estimators may perform worse than corresponding singly robust estimators (Kang and Schafer, 2007), although this was not observed in the simulation study presented in Section 4. Regardless of the estimator used, standard model-checking procedures should be employed in practice. A limitation of the data illustration in Section 5 was the assumption of no measurement error in covariates, which may not hold, e.g., for BMI. Finally, if the two markers investigated in this application are immunologically linked, then the causal estimands targeted by the two individual analyses may be poorly defined or difficult to interpret since changing one marker would necessarily change the other.
There are several possible extensions of the methods described in this paper. The proposed methods only account for continuous mismeasured exposures, but similar approaches could be developed to handle misclassification of categorical exposures. While this paper focuses on exposure measurement error, covariate measurement error is also common in practice and has been considered in causal analyses (Kuroki and Pearl, 2014; Lockwood and McCaffrey, 2015; Kyle et al., 2016; Hong, Rudolph, and Stuart, 2017); an approach which adjusts for both exposure and covariate measurement error would be a useful extension of the proposed methods. Methods could be developed that accommodate different measurement error models. Finally, methods that leverage non-parametric machine-learning approaches instead of parametric models could allow for more flexible inference.
Another possible future direction of research would be to combine g-estimation and CS methods to estimate parameters of a structural nested model where at least one exposure is measured with error. In addition, while this paper considers corrected score estimation methods, the conditional score estimation procedure (Carroll et al., 2006) is a related method which could be extended to a causal inference setting in similar ways. The conditional score also adjusts for measurement error without individual-level supplemental data and has advantages and disadvantages compared to CS methods in various settings. Finally, this paper focuses on point-exposures, but CS based approaches addressing measurement error for causal inference using longitudinal and survival data could also be developed.
Acknowledgements
The authors thank the Editor, Associate Editor, and two reviewers, as well as Kayla Kilpatrick, Shaina Mitchell, Sam Rosin, Bonnie Shook-Sa, and Jaffer Zaidi for helpful comments and suggestions. The authors also thank the investigators, staff, and participants of the HVTN 505 trial.
Research reported in this publication was supported by U.S. Public Health Service Grant AI068635, the National Institute of Allergy And Infectious Diseases of the National Institutes of Health (NIH) under Award Number R37 AI054165, and the National Institute of Environmental Health Sciences of the NIH under Award Number T32 ES007018. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Data Availability Statement
The data that support the findings of this study are available on Atlas at
https://atlas.scharp.org/cpas/project/HVTN%20Public%20Data/HVTN%20505/begin.view.
Supporting Information
Web Appendices, Tables, and Figures referenced in Sections 34, and 5 are available at the end of this document. All R code used in the simulations and application is available at the Biometrics website or in the R package mismex at https://github.com/brian-d-richardson/mismex.
Supporting Information for
July 12, 2024
A Large sample properties
In this appendix, the large sample properties of the proposed estimators discussed in Section 3 of the main paper are proven. In particular, the g-formula, IPW, and doubly-robust estimating functions $\boldsymbol{\Psi}_{GF-0},\boldsymbol{\Psi}_{IPW-0},\boldsymbol{\Psi}_{DR-0}$ which ignore measurement error are shown to be unbiased (i.e., have expected value equal to $\boldsymbol{0}$) and the components that involve $\boldsymbol{A}$ are shown to be conditionally unbiased given $\boldsymbol{A}$. Given these results, it follows that the three proposed corrected score estimators are consistent and asymptotically normal (Carroll et al., 2006).
A.1 G-formula CS estimator
Let $\boldsymbol{\Psi}^{(k)}$ denote the $k^{\rm th}$ row of an estimating function $\boldsymbol{\Psi}$. Consider the conditional expectation of the first row of the g-formula estimating function,
$\text{E}\left\{\boldsymbol{\Psi}_{GF-0}^{(1)}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\theta}_{GF})|\boldsymbol{A}\right\}$ $=\text{E}\left[\left\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\right\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})|\boldsymbol{A}\right]$
$=\text{E}\left(\text{E}\left[\left\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\right\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})|\boldsymbol{L},\boldsymbol{A}\right]|\boldsymbol{A}\right)$
$=\text{E}\left[\left\{\underbrace{\text{E}(Y|\boldsymbol{L},\boldsymbol{A})-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})}_{=0}\right\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})|\boldsymbol{A}\right]$
$=\boldsymbol{0},$
where the last inequality follows by the assumption of a correctly specified outcome model. The second row of $\boldsymbol{\Psi}_{GF-0}$, which does not involve $\boldsymbol{A}$, has marginal expectation
$\text{E}\left\{\boldsymbol{\Psi}_{GF-0}^{(2)}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\theta}_{GF})\right\}$ $=\text{E}\left\{\eta(\boldsymbol{a})-\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta})\right\}$
$=\eta(\boldsymbol{a})-\text{E}\left\{\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta})\right\}$
[…]
$=\eta(\boldsymbol{a})-\text{E}\left[\text{E}\{Y(\boldsymbol{a})|\boldsymbol{L},\boldsymbol{A}=\boldsymbol{a}\}\right]$ (causal consistency)
$=\eta(\boldsymbol{a})-\text{E}\left[\text{E}\{Y(\boldsymbol{a})|\boldsymbol{L}\}\right]$ (conditional exchangeability)
$=\eta(\boldsymbol{a})-\text{E}\{Y(\boldsymbol{a})\}$
$=\boldsymbol{0}.$
A.2 IPW CS estimator
For the IPW estimator, it is assumed that the parameters $\boldsymbol{\xi},\boldsymbol{\zeta}$ in the propensity models for $\boldsymbol{A}$ and $\boldsymbol{A}|\boldsymbol{L}$ are consistently estimated with the root of $\sum_{i=1}^{n}\widetilde{\boldsymbol{\Psi}}_{PS}(\boldsymbol{L}_{i},\boldsymbol{A}^{*}_{i},\boldsymbol{\Sigma}_{me},\boldsymbol{\xi},\boldsymbol{\zeta})$. Then it remains to show that the second row of $\boldsymbol{\Psi}_{IPW-0}$ is conditionally unbiased.
$\text{E}\left\{\boldsymbol{\Psi}_{IPW-0}^{(2)}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\theta}_{IPW})|\boldsymbol{A}=\boldsymbol{a}\right\}$
$=\text{E}\left[SW(\boldsymbol{L},\boldsymbol{A})\left\{Y-\eta(\boldsymbol{A};\boldsymbol{\gamma})\right\}\partial_{\boldsymbol{\gamma}}\eta(\boldsymbol{A};\boldsymbol{\gamma})|\boldsymbol{A}=\boldsymbol{a}\right]$
$=\text{E}\left[\frac{f_{\boldsymbol{A}}(\boldsymbol{A};\boldsymbol{\xi})}{f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L})}\left\{Y-\eta(\boldsymbol{A};\boldsymbol{\gamma})\right\}\partial_{\boldsymbol{\gamma}}\eta(\boldsymbol{A};\boldsymbol{\gamma})|\boldsymbol{A}=\boldsymbol{a}\right]$
$=f_{\boldsymbol{A}}(\boldsymbol{a};\boldsymbol{\xi})\partial_{\boldsymbol{\gamma}}\eta(\boldsymbol{a};\boldsymbol{\gamma})\text{E}\left\{\frac{Y(\boldsymbol{a})-\eta(\boldsymbol{a};\boldsymbol{\gamma})}{f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{a}|\boldsymbol{L})}|\boldsymbol{A}=\boldsymbol{a}\right\}$
[…]
$=f_{\boldsymbol{A}}(\boldsymbol{a};\boldsymbol{\xi})\partial_{\boldsymbol{\gamma}}\eta(\boldsymbol{a};\boldsymbol{\gamma})f_{\boldsymbol{A}}(\boldsymbol{a})^{-1}\left[\text{E}\{Y(\boldsymbol{a})\}-\eta(\boldsymbol{a};\boldsymbol{\gamma})\right]$
$=\boldsymbol{0}.$
Note that this conditional unbiasedness result does not rely on $f_{\boldsymbol{A}}(\boldsymbol{A};\boldsymbol{\xi})$ being the correct marginal density for $\boldsymbol{A}$, but it does rely on $f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L};\boldsymbol{\zeta})$ being the correct conditional density for $\boldsymbol{A}|\boldsymbol{L}$.
A.3 Doubly robust CS estimator
To show that the doubly robust estimator is in fact doubly robust, we show that $\boldsymbol{\Psi}_{DR-CS}$ is unbiased when either (i) the outcome model $\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})$ or (ii) the propensity model $f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A},\boldsymbol{L};\boldsymbol{\xi})$ is correctly specified. In the former case, we show that the second and third rows of $\boldsymbol{\Psi}_{DR-CS}$ are unbiased at the true values of $\boldsymbol{\beta},\eta(\boldsymbol{a})$ and at arbitrary values of $\boldsymbol{\xi},\boldsymbol{\zeta}$. In the latter case, we show that the first and third rows of $\boldsymbol{\Psi}_{DR-CS}$ are unbiased at the true values of $\boldsymbol{\xi},\boldsymbol{\zeta},\eta(\boldsymbol{a})$ and at the root $\boldsymbol{\beta}^{*}$ of the expected value of the second row. In both cases, the unbiasedness of the third row, which corresponds to the parameter of interest $\eta(\boldsymbol{a})$, implies consistent estimation of $\eta(\boldsymbol{a})$.
A.3.1 Correctly specified outcome model
Suppose the outcome model is correctly specified such that $\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})=\text{E}(Y|\boldsymbol{L},\boldsymbol{A})$ at the true value of $\boldsymbol{\beta}$, and the propensity model is possibly misspecified such that the denominator of $SW(\boldsymbol{L},\boldsymbol{A})$ may not equal the true density $f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L})$. Then the second row of $\boldsymbol{\Psi}_{DR-0}$ has conditional expectation
$\text{E}\left[SW(\boldsymbol{L},\boldsymbol{A})\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})|\boldsymbol{A}\right]$
$=\text{E}\left(\text{E}\left[SW(\boldsymbol{L},\boldsymbol{A})\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})|\boldsymbol{L},\boldsymbol{A}\right]|\boldsymbol{A}\right)$
$=\text{E}\left(SW(\boldsymbol{L},\boldsymbol{A})\underbrace{\{\text{E}(Y|\boldsymbol{L},\boldsymbol{A})-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})\}}_{=0}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})|\boldsymbol{A}\right)$
$=\boldsymbol{0}.$
The third row of $\boldsymbol{\Psi}_{DR-0}$ equals the second row of $\boldsymbol{\Psi}_{GF-0}$, which was shown in Web Appendix A.1 to be marginally unbiased when the outcome model is correctly specified..
A.3.2 Correctly specified propensity model
Now suppose the propensity model is correct but the outcome model may be incorrect. That is, $SW(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})=h(\boldsymbol{A})/f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L})$ for some function $h$, but the posited $\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta})$ may not equal the true $\text{E}(Y|\boldsymbol{L},\boldsymbol{A})$ for any value of $\boldsymbol{\beta}$. Let $\boldsymbol{\beta}^{*}=\{\beta^{*}_{0},\boldsymbol{\beta}^{*}_{l},\boldsymbol{\beta}^{*}_{a},\text{vec}(\boldsymbol{\beta}^{*}_{al})\}$ be the root of $\text{E}\{\boldsymbol{\Psi}_{DR-CS}^{(2)}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\beta})\}$. Then we claim that $\boldsymbol{\Psi}_{DR-0}$ is unbiased at $\boldsymbol{\theta}^{*}_{DR}=\{\boldsymbol{\beta}^{*},\eta(\boldsymbol{a};\boldsymbol{\gamma}),\boldsymbol{\xi},\boldsymbol{\zeta}\}$, where $\eta(\boldsymbol{a};\boldsymbol{\gamma})$ is the true MSM. The proof has two steps: (i) to show that $\boldsymbol{\beta}^{*}$ is the root of $\text{E}\left\{\boldsymbol{\Psi}_{DR-0}^{(2)}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\beta})|Y,\boldsymbol{L}\right\}$, and (ii) to show that this implies $\text{E}\{Y(\boldsymbol{a})\}=\text{E}\{\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta}^{*})\}$.
For the first step, we will use use equation (7.29) in Chapter 7 of Carroll et al. (2006) which says that, for a suitably smooth and integrable function $g(\boldsymbol{A})$,
$g(\boldsymbol{A})=\text{E}\left(\text{E}\left[\text{Re}\left\{g(\widetilde{\boldsymbol{A}})\right\}|Y,\boldsymbol{A}^{*},\boldsymbol{L}\right]|\boldsymbol{A}\right).$ (14)
Taking the expectation of both sides of (14) implies
$\text{E}\{g(\boldsymbol{A})\}$ $=\text{E}\left\{\text{E}\left(\text{E}\left[\text{Re}\left\{g(\widetilde{\boldsymbol{A}})\right\}|Y,\boldsymbol{A}^{*},\boldsymbol{L}\right]|\boldsymbol{A}\right)\right\}$
$=\text{E}\left(\text{E}\left[\text{Re}\left\{g(\widetilde{\boldsymbol{A}})\right\}|Y,\boldsymbol{A}^{*},\boldsymbol{L}\right]\right)$
$=\text{E}\left[\text{Re}\left\{g(\widetilde{\boldsymbol{A}})\right\}\right].$ (15)
By the definition of $\boldsymbol{\beta}^{*}$, and letting $g(\boldsymbol{A})=\boldsymbol{\Psi}_{DR-0}^{(2)}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\beta}^{*})$,
$\boldsymbol{0}$ $=\text{E}\left\{\boldsymbol{\Psi}_{DR-CS}^{(2)}(Y,\boldsymbol{L},\boldsymbol{A}^{*},\boldsymbol{\beta}^{*})\right\}$ (definition of $(\text{definition of }\beta^{*})$
[…]
$=\text{E}\left[\text{E}\left(\text{E}\left[\text{Re}\left\{\boldsymbol{\Psi}_{DR-0}^{(2)}(Y,\boldsymbol{L},\widetilde{\boldsymbol{A}},\boldsymbol{\beta}^{*})\right\}|Y,\boldsymbol{L},\boldsymbol{A}^{*}\right]|Y,\boldsymbol{L}\right)\right]$ (definition of $(\text{definition of }\boldsymbol{\Psi}_{DR-CS}^{(2)})$
[…]
The second step follows the proof in Zhang et al. (2016) of double robustness of this estimator in the absence of measurement error. Specifically, (A.3.2) implies
$\boldsymbol{0}$ $=\text{E}\left\{\boldsymbol{\Psi}_{DR-0}^{(2)}(Y,\boldsymbol{L},\boldsymbol{A},\boldsymbol{\beta}^{*})\right\}$
$=\text{E}\left[SW(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})\left\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta}^{*})\right\}\partial_{\boldsymbol{\beta}}\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta}^{*})\right],$
which in turn implies
$\boldsymbol{0}$ $=\text{E}\left[SW(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta}^{*})\}(1,\boldsymbol{A})^{T}\right]$
$=\text{E}\left(\text{E}\left[SW(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})\{Y-\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta}^{*})\}(1,\boldsymbol{A})^{T}|\boldsymbol{A}\right]\right)$
$=\text{E}\left(h(\boldsymbol{A})\left[\text{E}\left\{\frac{Y}{f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L})}|\boldsymbol{A}\right\}-\text{E}\left\{\frac{\mu(\boldsymbol{L},\boldsymbol{A};\boldsymbol{\beta}^{*})}{f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L})}|\boldsymbol{A}\right\}\right](1,\boldsymbol{A})^{T}\right).$ (17)
By arguments similar to those in Appendix A.2, it can be shown that
$h(\boldsymbol{A})E\left\{\frac{Y}{f_{\boldsymbol{A}|\boldsymbol{L}}(\boldsymbol{A}|\boldsymbol{L})}|\boldsymbol{A}\right\}=\eta(\boldsymbol{A};\boldsymbol{\gamma})$
and
[…]
$=\eta(\boldsymbol{A};\boldsymbol{\gamma}^{*}),$
where $\boldsymbol{\gamma}^{*}$ is obtained by replacing $\boldsymbol{\beta}$ for $\boldsymbol{\beta}^{*}$ in the expression for $\boldsymbol{\gamma}$. Then substituting these expressions into (A.3.2) gives
$\boldsymbol{0}$ $=\text{E}\left[\left\{\eta(\boldsymbol{A};\boldsymbol{\gamma})-\eta(\boldsymbol{A};\boldsymbol{\gamma}^{*})\right\}(1,\boldsymbol{A})^{T}\right],$
which implies that $\boldsymbol{\gamma}^{*}=\boldsymbol{\gamma}$. Thus, $\text{E}\{Y(\boldsymbol{a})\}=\text{E}\{\mu(\boldsymbol{L},\boldsymbol{a};\boldsymbol{\beta}^{*})\}$ and $\boldsymbol{\Psi}_{DR-CS}$ is unbiased at $\boldsymbol{\theta}^{*}_{DR}=\{\boldsymbol{\beta}^{*},\eta(\boldsymbol{a};\boldsymbol{\gamma}),\boldsymbol{\xi},\boldsymbol{\zeta}\}$.
B Closed form corrected score functions
In Section 3.5 of the main text, three examples of corrected score functions with known closed forms are mentioned. Here, a closed form for the IPW CS function is provided in the case where the PS models for $\boldsymbol{A}$ and $\boldsymbol{A}|\boldsymbol{L}$ are normal with identity links and where the MSM has a linear link. The same strategy can be used to find a closed form for the DR CS function with the same type of PS models, and where the outcome model has an identity link.
For the sake of simplicity, consider a univariate exposure $A$ subject to measurement error $\epsilon\sim\mathcal{N}(0,\sigma_{me}^{2})$, where $A|\boldsymbol{L}\sim\mathcal{N}(\mu_{\boldsymbol{L}},\delta^{2})$, $A\sim\mathcal{N}(\mu,\tau^{2})$, and the PS model parameters are known. This result can generalize to multidimensional $\boldsymbol{A}$ with parameters estimated using the root of $\boldsymbol{\Psi}_{PS}(\boldsymbol{L},\boldsymbol{A},\boldsymbol{\xi},\boldsymbol{\zeta})$. In this setting and in the absence of measurement error, the IPW estimating function is
$\boldsymbol{\Psi}_{IPW-0}(Y,\boldsymbol{L},A,\boldsymbol{\theta}_{IPW})=SW(\boldsymbol{L},A)\left\{Y-(\gamma_{0}-\gamma_{a}A)\right\}(1,A)^{T},$
where, using the normality of $A$ and $A|\boldsymbol{L}$,
$SW(\boldsymbol{L},A)$ $=\frac{\delta}{\tau}\exp\left\{\underbrace{\frac{1}{2}\left(\delta^{-2}-\tau^{-2}\right)}_{b_{1}}A^{2}+\underbrace{\left(\frac{\mu}{\tau^{2}}-\frac{\mu_{\boldsymbol{L}}}{\delta^{2}}\right)}_{b_{2}}A+\underbrace{\frac{1}{2}\left(\frac{\mu^{2}}{\tau^{2}}-\frac{\mu_{\boldsymbol{L}}^{2}}{\delta^{2}}\right)}_{b_{3}}\right\}$
$=\frac{\delta}{\tau}\exp\left(b_{1}A^{2}+b_{2}A+b_{3}\right).$
Then the IPW CS estimating function is $\boldsymbol{\Psi}_{IPW-CS}(Y,\boldsymbol{L},A^{*},\sigma^{2}_{me},\boldsymbol{\theta}_{IPW})$
$=\text{E}\left(\text{Re}\left[SW(\boldsymbol{L},\widetilde{A})\left\{Y-(\gamma_{0}+\gamma_{1}\widetilde{A})\right\}(1,\widetilde{A})^{T}\right]|Y,\boldsymbol{L},A^{*}\right)$
$=\text{E}\left\{\text{Re}\left(SW(\boldsymbol{L},A^{*}+i\widetilde{\epsilon})\left[Y-\{\gamma_{0}+\gamma_{1}(A^{*}+i\widetilde{\epsilon})\}\right](1,A^{*}+i\widetilde{\epsilon})^{T}\right)|Y,\boldsymbol{L},A^{*}\right\}.$ (18)
Note that the expectation in (B) is conditional on $Y,\boldsymbol{L},A^{*}$, and the only random component is $\widetilde{\epsilon}$. With this in mind, expressions in (B) can be written as
$SW(\boldsymbol{L},A^{*}+i\widetilde{\epsilon})$ $=\frac{\delta}{\tau}\exp\left\{\underbrace{b_{1}A^{*2}+b_{2}A^{*}+b_{3}}_{c_{1}}-\underbrace{b_{1}}_{c_{2}}\widetilde{\epsilon}^{2}+\underbrace{(2b_{1}+b_{2})}_{c_{3}}i\widetilde{\epsilon}\right\}$
$=\frac{\delta}{\tau}\exp\left(c_{1}-c_{2}\widetilde{\epsilon}^{2}+c_{3}i\widetilde{\epsilon}\right),$
and
$\left[Y-\{\gamma_{0}-(A^{*}+i\widetilde{\epsilon})\}\right]\begin{bmatrix}1\\ A^{*}+i\widetilde{\epsilon}\end{bmatrix}$ $=\begin{bmatrix}\overbrace{Y-\gamma_{0}-\gamma_{1}A^{*}}^{d_{1}}-\overbrace{\gamma_{2}}^{d_{2}}i\widetilde{\epsilon}\\ \underbrace{(Y-\gamma_{0})-\gamma_{1}A^{*2}}_{d_{3}}+\underbrace{\gamma_{1}}_{d_{4}}\widetilde{\epsilon}^{2}+\underbrace{(Y-\gamma_{0}-2\gamma_{1}A^{*})}_{d_{5}}i\widetilde{\epsilon}\end{bmatrix}$
$=\begin{bmatrix}d_{1}-d_{2}i\widetilde{\epsilon}\\ d_{3}+d_{4}\widetilde{\epsilon}^{2}+d_{5}i\widetilde{\epsilon}\end{bmatrix},$
where $c_{1},c_{2},c_{3},d_{1},d_{2},d_{3},d_{4},d_{5}$ are real constants with respect to the conditional expectation. Then (B) can be rewritten as
$\frac{\delta}{\tau}\text{E}\left[\text{Re}\left\{\exp\left(c_{1}-c_{2}\widetilde{\epsilon}^{2}+c_{3}i\widetilde{\epsilon}\right)\begin{bmatrix}d_{1}-d_{2}i\widetilde{\epsilon}\\ d_{3}+d_{4}\widetilde{\epsilon}^{2}+d_{5}i\widetilde{\epsilon}\end{bmatrix}\right\}\right].$ (19)
The expectation in (19) can be evaluated using Euler’s formula:
$\exp(i\alpha)$ $=\cos(\alpha)+i\sin(\alpha)\hskip 10.00002pt\forall\alpha\in\mathbb{R},$
and the characteristic function of $\widetilde{\epsilon}$:
$\text{E}\{\exp(i\widetilde{\epsilon})\}$ $=\exp(-\sigma_{me}^{2}/2),$
to show that $\boldsymbol{\Psi}_{IPW-CS}(Y,\boldsymbol{L},A^{*},\sigma^{2}_{me},\boldsymbol{\theta}_{IPW})$
$=\frac{\delta}{\tau}(1-2\sigma_{me}^{2}c_{2})^{-1/2}\exp\left\{c_{1}-\frac{c_{3}^{2}}{2(\sigma_{me}^{-2}-2c_{2})}\right\}\\ \times\begin{bmatrix}d_{1}+d_{2}c_{3}(\sigma_{me}^{-2}-2c_{2})^{-1}\\ d_{3}-d_{5}c_{3}(\sigma_{me}^{-2}-2c_{2})^{-1}+d_{4}(\sigma_{me}^{-2}-2c_{2})^{-1}\left(1-\frac{c_{3}^{2}}{\sigma_{me}^{-2}-2c_{2}}\right)\end{bmatrix}.$
C Two-phase sampling
C.1 Two-phase sampling method
Many studies (including the HVTN 505 trial) use a two-phase sampling design. Such a design is particularly useful when the outcome is easy to measure but the exposure of interest or some covariates are expensive or difficult to measure. Because each of the proposed methods above belongs to the estimating equation framework, it is straightforward to incorporate previously described methods for causal inference from studies with two-phase sampling. In this section, one such approach is demonstrated using a simulation study. In particular, for this simulation and the application section analysis, a simple inverse probability of sampling weights method is used (Wang et al., 2009).
The method is implemented by weighting each individual’s contribution to the estimating equations by the inverse probability of selection for the second-phase of the study (multiplying treatment weights by sampling weights for the IPW-CSM and DR-CSM estimators) and restricting the analysis to those selected. This method is well-suited for the subset of the HVTN 505 trial that is the focus of Section 5 of the main paper, particularly because all exposures of interest were measured in the second-phase sub-sample and no exposures were measured in the full sample.
C.2 Two-phase sampling simulations
The second simulation study described in Section 4 of the main paper was replicated, but under a two-phase sampling design and with the DR CS estimator instead of the IPW. In particular, a case-cohort design was used where the exposure was measured for a random sub-cohort as well as for every case. This was done for a sample size of $n=2000$ under three scenarios, with sub-cohorts of size $5\%$, $10\%$ and $25\%$. Since only the DR estimator was considered (and not the IPW estimator, which targets the MSM parameters $\boldsymbol{\gamma}$), the estimand of interest for this simulation study was $\eta(\boldsymbol{a})=\text{E}\{Y(\boldsymbol{a})\}$ for three values $\boldsymbol{a}_{1},\boldsymbol{a}_{2},\boldsymbol{a}_{3}\in\boldsymbol{\mathcal{A}}$. The results of 1000 simulation runs are presented in Web Table 1.
[…]
UC BC
Sub-cohort size Parameter Bias ESE ASE Cov ASE Cov
5 $\eta(\boldsymbol{a}_{1})$ 4.1 10.1 17.2 90.3 21.3 91.3
$\eta(\boldsymbol{a}_{2})$ 4.9 4.5 7.2 82.1 7.8 83.3
$\eta(\boldsymbol{a}_{3})$ 5.8 9.4 14.3 82.4 17.8 83.7
10 $\eta(\boldsymbol{a}_{1})$ 2.5 6.8 12.9 92.1 14.1 92.9
$\eta(\boldsymbol{a}_{2})$ 2.9 3.9 9.2 87.3 10.0 87.9
$\eta(\boldsymbol{a}_{3})$ 3.4 6.6 8.5 87.8 9.1 88.4
25 $\eta(\boldsymbol{a}_{1})$ 1.1 4.3 4.5 92.9 4.9 93.6
$\eta(\boldsymbol{a}_{2})$ 1.2 2.4 2.4 92.2 2.4 92.1
$\eta(\boldsymbol{a}_{3})$ 1.4 4.7 4.7 89.5 5.1 90.3
There was some bias and under-coverage when the sub-cohorts are smaller. This is likely due to a low effective sample size and tends to improve with increasing sample size. In addition, the estimator failed to converge in some settings. However, the DR CS estimator with sampling weights converged in all analyses presented in Section 5 of the main paper.
D Additional simulations
In this section, additional simulations are presented under two assumption violations: (i) near positivity violation and (ii) non-additive measurement error.
D.1 Near positivity violation
To evaluate the proposed g-formula CS method under a near positivity violation, the general structure of the first simulation study from Section 4 of the main paper was replicated almost exactly. A near positivity violation was created by changing how the treatment $A$ was generated from $\mathcal{N}(2+0.3L_{1}-0.5L_{2},0.6)$ to $\mathcal{N}(2+0.3L_{1}-0.5L_{2},0.35)$. Under this data generating process, which is illustrated in Web Figure 1, observations with $L_{2}=1$ were very unlikely to have exposure values greater than $3.5$.
Web Figure 1: Random samples from $A|\boldsymbol{L}$ from the near positivity violation simulations.
The results of the simulation study are presented in Web Figure 2 and Web Table 2, following the same format as Table 1 and Figure 2 in the main paper.
Web Figure 2: Estimated dose-response curve bias for the oracle, naive, and g-formula corrected score (CS) estimators under a near positivity violation. Bias refers to the average bias across 1000 simulated data sets for each method evaluated at each point on the horizontal axis, corresponding to setting the true exposure to $a\in[1,4]$.
Web Table 2: Results from the simulation study with a near positivity violation. UC, BC, Bias, ESE, ASE, and Cov as in Web Table 1. For 2 out of 1000 replicates with $n=800$, the g-formula CS estimator did not converge.
UC BC
n Method Bias ESE ASE Cov ASE Cov
800 Oracle G-Formula 0.3 3.2 3.3 95.1 3.4 95.7
800 Naive G-Formula -3.9 2.6 2.7 66.1 2.7 66.7
800 G-Formula CS 1.5 6.9 4.2 85.8 17.9 91.4
8000 Oracle G-Formula 0.0 1.1 1.0 93.8 1.0 93.8
8000 Naive G-Formula -4.1 0.9 0.8 0.2 0.8 0.2
8000 G-Formula CS 0.1 1.4 1.3 93.3 1.3 93.4
The results overall were similar to those in Table 1 and Figure 2 of the main paper, with slightly larger bias and ESE for each estimator. However, the proposed method still performed much better than the naive approach in this scenario.
D.2 Non-additive measurement error
The proposed methods were evaluated when treatment measurement error did not follow the classical additive model. In particular, the second simulation study from Section 4 of the main text was replicated, but the simulation of mismeasured treatment $A^{*}_{2}$ was changed such that it followed a multiplicative error model simulated as $A_{2}^{*}=A_{2}\epsilon_{me_{2}}$ where $\epsilon_{me_{2}}\sim\mathcal{N}(1,0.34)$. The proposed methods were still implemented (incorrectly) assuming additive measurement error with known measurement error covariance as specified in Section 4 of the main paper. The results are presented in Web Table 3.
Web Table 3: Results from the non-additive measurement error simulation study. UC, BC, Bias, ESE, ASE, and Cov as in Web Table 1.
UC BC
n Method Parameter Bias ESE ASE Cov ASE Cov
800 Oracle IPW $\gamma_{0}$ 0.0 3.5 3.2 92.9 3.2 93.2
$\gamma_{1}$0.4 8.0 7.3 91.6 7.4 92.2
$\gamma_{2}$ 0.0 8.1 7.3 91.7 7.4 92.7
$\gamma_{3}$ 0.1 7.2 6.9 93.5 7.0 93.8
Naive IPW $\gamma_{0}$ 0.4 2.8 2.7 93.1 2.7 93.6
$\gamma_{1}$7.2 5.6 5.4 73.4 5.5 73.9
$\gamma_{2}$5.4 6.8 5.8 82.4 6.0 84.2
$\gamma_{3}$ 0.2 6.7 6.5 94.5 6.5 94.5
Corrected IPW $\gamma_{0}$0.4 4.9 4.7 91.0 4.9 92.1
$\gamma_{1}$ 0.3 14.1 14.2 92.4 14.7 93.6
$\gamma_{2}$2.0 14.3 12.2 86.5 13.1 89.7
$\gamma_{3}$ 0.4 9.1 9.1 94.1 9.2 94.2
8000 Oracle IPW $\gamma_{0}$ 0.0 1.2 1.1 95.4 1.1 95.5
$\gamma_{1}$0.1 2.8 2.7 95.2 2.7 95.4
$\gamma_{2}$ 0.2 2.8 2.6 94.2 2.6 94.2
$\gamma_{3}$ 0.0 2.3 2.4 95.0 2.4 94.9
Naive IPW $\gamma_{0}$ 0.4 0.9 0.9 92.0 0.9 92.1
$\gamma_{1}$7.1 1.8 1.9 3.9 1.9 3.9
$\gamma_{2}$5.6 2.8 2.4 34.7 2.4 35.9
$\gamma_{3}$ 0.1 2.2 2.2 94.3 2.2 94.3
Corrected IPW $\gamma_{0}$0.7 4.9 9.0 90.9 8.9 93.6
$\gamma_{1}$ 0.8 20.6 36.8 93.6 35.6 94.5
$\gamma_{2}$3.5 8.8 6.8 84.8 7.5 88.6
$\gamma_{3}$ 0.3 4.8 7.4 93.3 7.3 93.7
The proposed IPW CS method continued to perform well for the MSM coefficients corresponding to treatments $A_{1}$ and $A_{3}$, for which the additivity assumption holds, but exhibited bias and undercoverage for the $A_{2}$ coefficient. However, the proposed CS estimator still outperformed the naive estimator in this setting.
D.3 Lower sample size
Finally, the proposed methods were evaluated under smaller sample sizes than that considered in the main paper. In particular, each of the three simulation studies was repeated with $n=400$. The replication of the first simulation study is presented in Web Table 4 and focuses on the point where $a=3$ rather than the entire dose-response curve. Likewise the replications of the second and third simulation studies are presented in Web Tables 5 and 6.
Web Table 4: Results from the first simulation study with sample size $n=400$. UC, BC, Bias, ESE, ASE, and Cov as in Web Table 1.
UC BC
n Method Bias ESE ASE Cov ASE Cov
400 Oracle G-Formula 0.4 4.0 4.0 93.9 4.2 94.5
Naive G-Formula -2.7 3.5 3.5 86.3 3.6 87.0
G-Formula CS 1.5 7.9 4.8 85.0 10.1 89.7
Web Table 5: Results from the second simulation study with sample size $n=400$. UC, BC, Bias, ESE, ASE, and Cov as in Web Table 1.
UC BC
n Method Parameter Bias ESE ASE Cov ASE Cov
400 Oracle IPW $\gamma_{0}$ 0.2 4.6 4.3 91.6 4.3 92.4
$\gamma_{1}$0.4 10.2 9.7 92.4 9.9 92.9
$\gamma_{2}$ 0.0 10.8 9.6 91.0 9.9 91.7
$\gamma_{3}$0.4 10.1 9.4 93.1 9.5 93.2
Naive IPW $\gamma_{0}$ 0.7 3.9 3.7 92.8 3.7 93.3
$\gamma_{1}$7.5 7.9 7.4 81.7 7.5 82.7
$\gamma_{2}$5.1 7.6 7.4 88.5 7.5 88.8
$\gamma_{3}$0.2 9.4 9.0 93.5 9.1 93.6
IPW CS $\gamma_{0}$ 0.2 7.5 9.2 89.5 9.6 90.4
$\gamma_{1}$ 0.2 16.1 16.3 89.7 17.0 90.8
$\gamma_{2}$ 1.2 22.6 28.3 90.3 29.8 92.0
$\gamma_{3}$0.2 11.5 11.2 92.7 11.5 92.9
Web Table 6: Results from the third simulation study with sample size $n=400$. UC, BC, Bias, ESE, ASE, and Cov as in Web Table 1.
UC BC
Correct Specifications Method Bias ESE ASE Cov ASE Cov
PS and OR DR CS 0.9 4.5 4.0 91.7 4.4 92.7
G-Formula CS 0.2 3.6 3.6 94.9 3.6 95.0
IPW CS -0.2 10.2 6.5 86.0 7.5 89.3
PS DR CS 0.3 7.1 5.5 89.1 6.0 90.8
G-Formula CS -6.8 4.0 4.0 60.9 4.1 61.9
IPW CS -0.2 10.2 6.5 86.0 7.5 89.3
OR DR CS 0.4 3.9 3.7 93.2 3.8 93.4
G-Formula CS 0.2 3.6 3.6 94.9 3.6 95.0
IPW CS -6.8 5.2 4.3 59.4 4.5 61.0
DR CS -6.4 4.7 4.1 61.8 4.2 62.9
G-Formula CS -6.8 4.0 4.0 60.9 4.1 61.9
IPW CS -6.8 5.2 4.3 59.4 4.5 61.0
Each of these simulations yielded similar results regarding bias of the estimators, compared to the simulations in the main text with larger sample size. However, in some cases, especially the g-formula simulations (Web Table 4), the bias of the proposed estimator and its corresponding uncorrected (UC) and bias-corrected (BC) variance estimators increased with the smaller sample size, resulting in undercoverage of the UC and BC CIs.
E Model diagnostics for the application section
Standard model diagnostics were used to evaluate both propensity model specifications in Section 5 of the main text. Diagnostics for the ADCP exposure are presented in Web Figure 3, which largely indicate a good model fit. Diagnostics for the RII exposure are presented in Web Figure 4. These diagnostics indicate two outliers for which the model had poor predictive performance.
Web Figure 3: Model diagnostics for the ADCP propensity model.
Web Figure 4: Model diagnostics for the RII propensity model.
To assess the fit of the outcome models, Chi-square goodness of fit tests were performed, yielding a p-value of 0.56 for the model including ADCP and a p-value of 0.33 for the model including RII. Based on the propensity model diagnostics, a supplemental analysis was performed with two outliers removed for the RII models (Web Figure 5). The diagnostics for this analysis no longer indicated poor propensity model fit (Web Figure 6) and yielded an outcome model Chi-square goodness of fit test p-value of 0.85. However, removing the outliers resulted in wider confidence intervals for lower levels of the exposure than in the analysis presented in Section 5 of the main paper.
Web Figure 5: Supplemental analysis removing two outliers in the RII models.
Web Figure 6: Model diagnostics for the RII propensity model in the supplemental analysis with two outliers removed.